OpenStudy (anonymous):

{}_)(*&^%^$@!$$# Help with a question I want someone to see if Im right???? ++!)I@!_!_)@((@_!(@(_!(((@"{}+(*^@#%&*())

OpenStudy (anonymous):

Line AB contains (0, 4) and (1, 6) Line CD contains points (2, 10) and (−1, 4). Lines AB and CD are A) parallel because the slopes are the same B) perpendicular because the slopes are the same C) parallel because the product of the slopes is −1 D) perpendicular because the product of the slopes is −1

OpenStudy (triciaal):

what was your answer?

OpenStudy (anonymous):

B

OpenStudy (triciaal):

no cannot be b or c perpendicular lines the product of the slope =-1 parallel lines the slopes =

OpenStudy (triciaal):

find the slope of each line

OpenStudy (anonymous):

6/3 10/3

OpenStudy (triciaal):

check your work and reduce to the lowest term

OpenStudy (triciaal):

slope 2

OpenStudy (anonymous):

can you show me how the slope is two?

OpenStudy (triciaal):

6/3 = 2

OpenStudy (anonymous):

ohhh ok

OpenStudy (triciaal):

6-4/1-0 =2/1 =2

OpenStudy (triciaal):

since the slopes are the same then they are parallel lines answer a

OpenStudy (anonymous):

ohh ok I understand now you think you can check a few more answers for me to see if im right?

OpenStudy (anonymous):

Line CD contains points A (4, −7) and B (4, 8). The slope of line CD is A) zero B) undefined <---- My answer C) positive D) negative

OpenStudy (triciaal):

8--7/4-4 agree

OpenStudy (anonymous):

Yay ! So its B???

OpenStudy (anonymous):

The equation of line AB is (y−3) = 5 (x − 4). What is the slope of a line perpendicular to line AB? A) −5 B) 5 C) -1/5 <---- my answer D) 1/5

OpenStudy (triciaal):

great!

OpenStudy (anonymous):

The equation of line WX is y = −2x − 5. Write an equation of a line perpendicular to line WX in slope-intercept form that contains point (−1, −2). A) y = 1/2x + 3/2 B) y = - 1/2x + 3/2 C) y = 1/2x − 3/2 D) y = − 1/2x − 3/2 <------ my answer

OpenStudy (triciaal):

-2 m =-1 m = -1/-2 = 1/2 point (-2, -1) -2 = 1/2 (-1) + b b = -2 + 1/2 = -3/2 y = 1/2 (x) -3/2

OpenStudy (triciaal):

ans c

OpenStudy (anonymous):

so its C?

OpenStudy (anonymous):

Ok last one

OpenStudy (anonymous):

The equation of line CD is y = −2x − 2. Write an equation of a line parallel to line CD in slope-intercept form that contains point (4, 5). A) y = −2x + 13 <---- My answer B) y = negative 1/2x + 7 C) y = negative 1/2x + 3 D) y = − 2x − 3

OpenStudy (triciaal):

show your work

OpenStudy (anonymous):

ok give me a min

OpenStudy (triciaal):

yes its correct!!